OpenStudy (anonymous):

A function is even if and only if f(-x) = f(x). A function is odd if and only if f(-x) = - f(x). Using these definitions together with the definition of the product of two functions, show that the product of two odd functions with the same domain is an even function.

OpenStudy (bahrom7893):

Product of two odd functions is: f(-x) * f(-x) = -f(x) * (-f(x)) = +g(x)

OpenStudy (anonymous):

So given that: \[f(-x) = f(x)\]\[g(-x) = -g(x)\] What is: \[f(-x)\cdot g(-x) ?\]

OpenStudy (anonymous):

thanks :)